OpenStudy (anonymous):

Simplify i36 A) 1 B) -1 C) -i D) i

OpenStudy (anonymous):

What's i?

OpenStudy (anonymous):

not sure, the question just tells me to simplify i36 and my lesson doesnt give anything on it

OpenStudy (anonymous):

A) 1

OpenStudy (anonymous):

Can you explain?

OpenStudy (2bornot2b):

\[i^{36}=(i^4)^9=1^9=1\]

OpenStudy (phi):

You must be i^36 (i to the 36 power) where i=sqrt(-1) sqrt(-1)*sqrt(-1)= -1 and -1*-1=1 so sqrt(-1)^4 (i^4) = +1 and (i^4)^9 = 1

OpenStudy (anonymous):

oh, I didn't take it in that sense.....

OpenStudy (anonymous):

I never guessed it would involve exponents.

hero (hero):

I like 2bornot2b's approach